# Multigroup Radiation Diffusion on a Moving Mesh: Implementation in rich and Application to Tidal Disruption Events

### v1

@@ Abstract
Radiation-hydrodynamics (RHD) determines the bulk evolution and observable emission in a wide variety of high-energy astrophysical phenomena. Due to their complexity, RHD problems must usually be studied through numerical simulation. We have extended the publicly available rich code, which previously solved the equations of RHD in the limit of grey flux-limited diffusion (FLD), to operate with a multigroup FLD solver. rich is a semi-Lagrangian code that solves the equations of RHD on an unstructured moving mesh, and is the first multigroup RHD moving mesh code, making it uniquely applicable to problems with extreme dynamic range and dynamically important radiation forces. We validate our multigroup module against multiple analytic benchmarks, including a novel test of the RHD Doppler term. The computational efficiency of the code is aided by a novel scheme to accelerate convergence in optically thick cells. Finally, we apply multigroup rich in a pilot study of a stellar tidal disruption event (TDE), using a $10^{4}M_{\odot}$ intermediate-mass black hole. Our simulations self-consistently produce a bright early-time X-ray flash prior to peak optical/UV light, in qualitative agreement with post-processing of (grey) rich simulations of supermassive black hole TDEs, as well as X-ray observations of the TDE AT 2022dsb.
@@ IV.2.2 Convergence Acceleration
where we typically set $N_{\tau}=100$, which limits the local equilibration time to be $\Delta t/N_{\tau}$ while leaving the diffusive transport coefficient $\kappa_{R,g}$ unchanged. In our tests, this reduces the iteration count substantially with negligible impact on the physical solution.
@@ V.1 Marshak Boundary Condition
The Marshak boundary condition describes the incoming flux into a medium, which is coupled to a heat bath at LTE with temperature $T_{bath}$. The boundary condition is given by (Krief and McClarren, 2024a, b; Derei et al., 2024)
@@ V.2 Grey Diffusion
In the optically thick limit ($\lambda_{g}\to 1/3$), and taking $\kappa_{P,g}$ and $\kappa_{R,g}$ to be constant, we can recover the grey radiation-diffusion equation by summing the multigroup equations over the energy groups
@@ V.2.1 Non-Equilibrium Nonlinear Marshak Waves
Figure 1: Comparison of the radiation temperature and material temperature at t=1nst=1\ $\text${ns} between the rich simulation and the analytic solution for Marshak problems 1 (left) and 2 (right). In problem 2 the absorption coefficient is large enough to enforce LTE, so $T_{\rm rad}\simeq T_{\rm mat}$.
@@ V.2.2 Mach 2 Radiative Shock
The simulation is performed using a uniform grid of 256 cells between x=1000x=-1000 cm and $x=2000$ cm, with $5$ energy groups between $10^{-3}\text{eV}$ and $1\text{eV}$ spaced logarithmically. We evolve the system until $t=0.02$ seconds, by then the system has evolved to the steady state solution. Figure 3 shows the gas temperature and radiation energy density profiles along with the semi‑analytic solution.
@@ V.3.1 Densmore Multigroup tests
Figure 4: Comparison of the gas temperature at $t=1$ ns between rich and the data taken from Densmore et al. (2012) for the cases $\kappa_{0}=10\;\text{keV}^{3.5}$ (left) and $\kappa_{0}=100\;\text{keV}^{3.5}$ (right)
@@
Figure 5: Comparison of the gas temperature at $t=1$ ns between rich and the data taken from Densmore et al. (2012) for the medium opacity case $\kappa_{0}=1000\;\text{keV}^{3.5}$ (left) and the step function opacity as presented in Eq. V.25 (right).
@@ V.3.3 Convergence Acceleration
The improved convergence method described in section IV.2.2 is evaluated using an astrophysically motivated benchmark. Specifically, we model a Marshak wave under physical conditions characteristic of Tidal Disruption Event simulations (Steinberg and Stone, 2024). The setup consists of a 1D domain with a Marshak boundary condition imposed at one end, where a bath temperature of $T_{bath}=10^{5}\;\text{K}$ drives radiation into cold gas of density $\rho=10^{-7}\;\text{g/cc}$. The gas follows an ideal-gas equation of state with adiabatic index $\gamma=5/3$ and a heat capacity of $C_{v}=10^{8}\rho\;\text{erg}\;\text{cm}^{-3}\;\text{K}^{-1}$. For the opacity we assume a free-free opacity described by $\kappa_{P,\nu}=\kappa_{R,\nu}=3.7\cdot 10^{8}n^{2}\nu^{-3}T^{-1/2}\cdot(1-\exp{(-h\nu/k_{B}T)})\;\text{cm}^{-1}$, assuming the gas is completely ionized at all times and all of the quantities are in CGS units. We evolve the system until a time of $t=10^{3}\;\text{s}$ on a uniform grid $x\in(0,\;10^{11})\;\text{cm}$ divided uniformly into 256 cells once with our convergence acceleration using $N_{\tau}=100$ and once without it. Figure 7 presents the gas temperature and the radiation temperatures in two energy groups as functions of position for both simulations. The profiles are nearly identical, confirming that the convergence-acceleration method does not alter the physical solution. However, the computational cost differs substantially: the run without acceleration required 923 s, whereas the accelerated run completed in only 82 s.
Figure 7: Temperature profiles of the gas (left) and of the radiation (right) in two energy groups (14 eV and 94 eV) for simulations conducted with and without the convergence-acceleration scheme. The overlapping curves confirm that the method maintains solution accuracy. The accelerated run reduces the computational cost by more than an order of magnitude, completing in 82 s compared to 923 s.
@@ V.4 Tidal Disruption Event
To demonstrate the astrophysical applicability of the new multigroup radiation–diffusion module, we employ it in a full, self-consistent tidal disruption event (TDE) simulation. We simulate the parabolic approach of a star with mass $M_{\star}=0.5M_{\odot}$ and radius $R_{\star}=0.47R_{\odot}$ towards a $M_{\bullet}=10^{4}M_{\odot}$ black hole. The star is quickly torn apart by the black hole’s tidal field, and its deconfined debris flies out to radii $\sim R_{\star}(M_{\bullet}/M_{\star})^{2/3}$ before the most bound material turns around at its apocenter. As stellar debris returns to pericenter, it begins dissipating orbital energy into heat through shocks. Eventually, a subset of this dissipated energy is radiated, producing the luminous nuclear flares seen today by time-domain observers (van Velzen et al., 2020). The details of shock dissipation, debris circularization, and radiative emission all remain hotly debated at present (Hayasaki et al., 2013; Piran et al., 2015; Shiokawa et al., 2015; Bonnerot and Stone, 2021; Bonnerot et al., 2021; Andalman et al., 2022; Steinberg and Stone, 2024; Huang et al., 2024; Andalman et al., 2025), and this proof-of-principle simulation does not aim at resolving major open questions in TDE physics. Because the simulation has the exact astrophysical parameters as the grey-rich simulations of (Martire et al., 2025), we generally expect a similar result: much weaker dissipation and slower circularization than what is seen in supermassive black hole TDEs; emission powered by dissipation at the nozzle shock that is reprocessed through a radiation-driven wind; near-Eddington peak luminosities. The main advantage of this work relative to past end-to-end TDE simulations is the use of multigroup (rather than grey) radiation transfer, which allows us to self-consistently predict light curve evolution in different frequency bands.
@@
Figure 8: Mass density slice in the orbital plane at $t=0.42\ t_{\rm fb}\approx 1\ \text{day}$ after the star was disrupted. The stellar debris has begun falling back, forming an accretion flow around the black hole. Gas densities are labeled according to the color bar in the figure, and are highest in the debris stream.
We extract the luminosity in each energy group along a given viewing angle $\vec{\Omega}$ (the angular grid is generated using HEALPix with $N_{\rm side}=8$; Górski et al. 2005)
@@
Figure 9: Mollweide sky maps of the Optical+UV luminosity (left) and the Soft+Hard X-ray band luminosities (right), at $t=0.4\ t_{fb}\approx 1\ \text{day}$. At this early stage of evolution, very little luminosity is being produced, but hints of later trends have already emerged: optical+UV emission is primarily emitted in the orbital plane where the reprocessing column density is highest, while X-ray emission is suppressed in the orbital plane for the same reason.
@@
In Figs. 911 we show isotropic equivalent Mollweide skymaps of band luminosities at t=0.40, 0.79,t=0.40,\,0.79, and $2.08\,t_{\rm fb}$, respectively. Figure 12 also presents ratio maps (Optical+UV/X-ray and Optical+UV/He) at $t=0.79\,t_{\rm fb}$. Near peak luminosity, the Optical+UV emission is strongest along the disrupted star’s orbital plane, consistent with reprocessing of an ionizing continuum in dense debris, whereas the X-ray bands peak at higher inclinations off the plane, where the reprocessing column density is lower (see especially Figs. 1012). At $t=2.08\,t_{\rm fb}$ a localized X-ray bright spot appears (Fig. 11), density and temperature slices indicate it arises from hot gas breaking out of the inner disk (Figs. 1415).
We further quantify the observable features of this simulated TDE by plotting sky-averaged multiband light curves, in Fig. 13. The sky-averaged light curves exhibit a clear X‑ray flash, peaking at $\approx 2.5$ days after disruption (Fig. 13). This result appears consistent with earlier predictions of similar X-ray flashes in supermassive black hole TDEs from Steinberg and Stone (2024), which appear to have been observed in the (supermassive) TDE AT 2022dsb (Malyali et al., 2024). In contrast to Steinberg and Stone (2024), we have made self-consistent predictions for the time-dependent X-ray luminosity in this paper using our novel multigroup FLD transfer module; with grey simulations, spectral predictions can only be made in post-processing, which may not be self-consistent in situations like TDE circularization where (frequency-dependent) radiation pressure dominates over purely hydrodynamical forces.
@@
Figure 13: Light curves for the different energy groups. Most radiation emerges in UV wavelengths; primarily in the unobservable EUV bands (H and He). UV bands (including Swift UVW2) reach peaks or plateaus near $t\approx 1t_{\rm fb}$, while optical wavelengths monotonically rise for the duration of the simulation. The earliest X-ray flash (at $t\approx 0.6~{\rm day}$) is likely an unphysical artifact of under-resolving the tip of the stream as it comes back to pericenter (Martire et al., 2025), but the broader X-ray flare from $\approx 1$ to $\approx 2.5$ days is more likely physical, and agrees qualitatively with both the simulation of Steinberg and Stone (2024) and observations of the TDE AT 2022dsb (Malyali et al., 2024). This paper’s multigroup simulations are the first self-consistent prediction of an early X-ray flash in an end-to-end simulation, suggesting that these flares may offer observational insight into both circularization mechanisms and viewing angles.
@@ VI Conclusions
Our implementation of multigroup FLD into rich expands the capabilities of moving-mesh RHD at the cost of greater complexity and computational expense. The multi-step RHD algorithm is described in more detail in §IV, but in brief, the code first updates the hydrodynamic variables, then moves the Voronoi mesh to follow the bulk fluid flow, and then updates the FLD radiation equation [^2] This overview is slightly oversimplified; for example, radiation advection occurs during the hydrodynamics sub-step.. Finally, radiation forces update the gas momentum and energy variables. To mitigate the greater computational cost of performing the RHD in a multigroup way, we have implemented a novel scheme to accelerate convergence in optically thick cells, with minimal loss of accuracy (as can be seen in Fig. 7).
@@
As a final test for multigroup rich, we have gone beyond test problems with analytic solutions and have run the first end-to-end simulation of a TDE in multigroup RHD. Our simulation follows the disruption of a lower main sequence star by a $M=10^{4}M_{\odot}$ IMBH, using astrophysical parameters identical to the grey RHD simulations recently presented in Martire et al. (2025). This simulation produced multi-band light curves across 10 different frequency groups, and runs for a little over 2 fallback times. Although we will defer more detailed explorations of multigroup TDE simulations for future work (Giron et al. in prep), this proof-of-principle simulation is already quite interesting. The bulk hydrodynamics and bolometric light curve appear broadly consistent with the slow circularization seen in Martire et al. (2025), but the emergent radiation field exhibits a complex angular dependence and frequency evolution. Optical and near-UV bands rise monotonically but never make up more than $\sim 1\%$ of the bolometric emission, which is released primarily in the unobservable EUV (although we note that the far-UV Swift UVW2 band does eventually reach $\sim 10\%$ of the bolometric luminosity).

### v2

@@ Abstract
Radiation-hydrodynamics (RHD) determines the bulk evolution and observable emission in a wide variety of high-energy astrophysical phenomena. Due to their complexity, RHD problems must usually be studied through numerical simulation. We have extended the publicly available rich code, which previously solved the equations of RHD in the limit of grey flux-limited diffusion (FLD), to operate with a multigroup FLD solver. rich is a semi-Lagrangian code that solves the equations of RHD on an unstructured moving mesh, and is the first multigroup RHD moving mesh code, making it uniquely applicable to problems with extreme dynamic range and dynamically important radiation forces. We validate our multigroup module against multiple analytic benchmarks, including a novel test of the RHD Doppler term. The computational efficiency of the code is aided by a novel scheme to accelerate convergence in optically thick cells by limiting the absorption coefficients. Finally, we apply multigroup rich in a pilot three dimensional study of a stellar tidal disruption event (TDE), using a $10^{4}M_{\odot}$ intermediate-mass black hole. Our simulations self-consistently produce a bright early-time X-ray flash prior to peak optical/UV light, in qualitative agreement with post-processing of (grey) rich simulations of supermassive black hole TDEs, as well as X-ray observations of the TDE AT 2022dsb.
@@ IV.2.2 Convergence Acceleration
where $N_{\tau}$ is the optical depth threshold, which we typically set $N_{\tau}=100$, which limits the local equilibration time to be $\Delta t/N_{\tau}$ while leaving the diffusive transport coefficient $\kappa_{R,g}$ unchanged. In our tests, this reduces the iteration count substantially with negligible impact on the physical solution.
@@ V.1 Marshak Boundary Condition
The Marshak boundary condition describes the incoming flux into a medium, which is coupled to a heat bath at LTE with temperature $T_{bath}$. LTE means the radiation has a Planck spectrum with the same temperature as the material temperature. The boundary condition is given by (Krief and McClarren, 2024a, b; Derei et al., 2024)
@@ V.2 Grey Diffusion
In the optically thick limit (i.e. small mean free path and gradient scale, $R_{g}\rightarrow 0$ and $\lambda_{g}\to 1/3$), when taking $\kappa_{P,g}$ and $\kappa_{R,g}$ to be constant, we can recover the grey radiation-diffusion equation by summing the multigroup equations over the energy groups
@@ V.2.1 Non-Equilibrium Nonlinear Marshak Waves
Figure 1: Comparison of the radiation temperature $T_{\rm rad}$ and material temperature $T_{\rm mat}$ at t=1nst=1\ $\text${ns} between the rich simulation and the analytic solution for Marshak problems 1 (left) and 2 (right). In problem 2 the absorption coefficient is large enough to enforce LTE, so $T_{\rm rad}\simeq T_{\rm mat}$.
@@ V.2.2 Mach 2 Radiative Shock
Where $c_{v}=\frac{\partial u_{m}}{\partial T}$ is the specific heat capacity. The simulation is performed using a uniform grid of 256 cells between x=1000x=-1000 cm and $x=2000$ cm, with $5$ energy groups between $10^{-3}\text{eV}$ and $1\text{eV}$ spaced logarithmically. We evolve the system until $t=0.02$ seconds, by then the system has evolved to the steady state solution. Figure 3 shows the gas temperature and radiation energy density profiles along with the semi‑analytic solution  [^1] For the Mach 2 radiative problem, we used the analytical solver by Ulrich Noebauer, available in the public-astro-tools repository at https://github.com/unoebauer/public-astro-tools/tree/master. .
@@ V.3.1 Densmore Multigroup tests
Figure 4: Comparison of the gas temperature at $t=1$ ns between rich and the Monte-Carlo simulation results taken from Densmore et al. (2012) for the cases $\kappa_{0}=10\;\text{keV}^{3.5}$ (left) and $\kappa_{0}=100\;\text{keV}^{3.5}$ (right)
@@
Figure 5: Comparison of the gas temperature at $t=1$ ns between rich and the Monte-Carlo simulation results taken from Densmore et al. (2012) for the medium opacity case $\kappa_{0}=1000\;\text{keV}^{3.5}$ (left) and the step function opacity as presented in Eq. V.25 (right).
@@ V.3.3 Convergence Acceleration
The improved convergence method described in section IV.2.2 is evaluated using an astrophysically motivated benchmark. Specifically, we model a Marshak wave under physical conditions characteristic of Tidal Disruption Event simulations (Steinberg and Stone, 2024). The setup consists of a 1D domain with a Marshak boundary condition imposed at one end, where a bath temperature of $T_{bath}=10^{5}\;\text{K}$ drives radiation into cold gas of density $\rho=10^{-7}\;\text{g/cc}$. The gas follows an ideal-gas equation of state with adiabatic index $\gamma=5/3$ and a heat capacity of $C_{v}=10^{8}\rho\;\text{erg}\;\text{cm}^{-3}\;\text{K}^{-1}$. For the opacity we assume a free-free opacity described by $\kappa_{P,\nu}=\kappa_{R,\nu}=3.7\cdot 10^{8}n^{2}\nu^{-3}T^{-1/2}\cdot(1-\exp{(-h\nu/k_{B}T)})\;\text{cm}^{-1}$, assuming the gas is completely ionized at all times and all of the quantities are in CGS units. We evolve the system until a time of $t=10^{3}\;\text{s}$ on a uniform grid $x\in(0,\;10^{11})\;\text{cm}$ divided uniformly into 256 cells once with our convergence acceleration using $N_{\tau}=100$ and once without it. Figure 7 presents the gas temperature and the radiation temperatures of two energy groups, $14\ \rm eV$ and $94\ \rm eV$ as functions of position for both simulations. The profiles are nearly identical, confirming that the convergence-acceleration method does not alter the physical solution. However, the computational cost differs substantially: the run without acceleration required 923 s, whereas the accelerated run completed in only 82 s.
Figure 7: Temperature profiles of the gas (left) and of the radiation (right) in two energy groups (14 eV and 94 eV) for simulations conducted with and without the limiting of the absorption coefficients. The overlapping curves confirm that the method maintains solution accuracy. The accelerated run reduces the computational cost by more than an order of magnitude, completing in 82 s compared to 923 s.
@@ V.4 Tidal Disruption Event
To demonstrate the astrophysical applicability of the new multigroup radiation–diffusion module, we employ it in a full, three dimensional, self-consistent tidal disruption event (TDE) simulation. We simulate the parabolic approach of a star with mass $M_{\star}=0.5M_{\odot}$ and radius $R_{\star}=0.47R_{\odot}$ towards a $M_{\bullet}=10^{4}M_{\odot}$ black hole. The star is quickly torn apart by the black hole’s tidal field, and its deconfined debris flies out to radii $\sim R_{\star}(M_{\bullet}/M_{\star})^{2/3}$ before the most bound material turns around at its apocenter. As stellar debris returns to pericenter, it begins dissipating orbital energy into heat through shocks. Eventually, a subset of this dissipated energy is radiated, producing the luminous nuclear flares seen today by time-domain observers (van Velzen et al., 2020). The details of shock dissipation, debris circularization, and radiative emission all remain hotly debated at present (Hayasaki et al., 2013; Piran et al., 2015; Shiokawa et al., 2015; Bonnerot and Stone, 2021; Bonnerot et al., 2021; Andalman et al., 2022; Steinberg and Stone, 2024; Huang et al., 2024; Andalman et al., 2025), and this proof-of-principle simulation does not aim at resolving major open questions in TDE physics. Because the simulation has the exact astrophysical parameters as the grey-rich simulations of (Martire et al., 2025), we generally expect a similar result: much weaker dissipation and slower circularization than what is seen in supermassive black hole TDEs; emission powered by dissipation at the nozzle shock that is reprocessed through a radiation-driven wind; near-Eddington peak luminosities. The main advantage of this work relative to past end-to-end TDE simulations is the use of multigroup (rather than grey) radiation transfer, which allows us to self-consistently predict light curve evolution in different frequency bands.
@@
Figure 8: Mass density slice of the TDE simulation in the orbital plane at $t=0.42\ t_{\rm fb}\approx 1\ \text{day}$ after the star was disrupted. The stellar debris has begun falling back, forming an accretion flow around the black hole. Gas densities are labeled according to the color bar in the figure, and are highest in the debris stream. The black hole is placed at the origin, $X,Y$ are in $R_{\odot}$.
We extract the luminosity in each energy group along a given viewing angle $\vec{\Omega}$ (the angular grid is generated using HEALPix with $N_{\rm side}=8$, Górski et al. 2005)
@@
Figure 9: Mollweide sky maps of the Optical+UV luminosity (left) and the Soft+Hard X-ray band luminosities (right), at $t=0.4\ t_{fb}\approx 1\ \text{day}$. At this early stage of evolution, very little luminosity is being produced, but hints of later trends have already emerged: optical+UV emission is primarily emitted in the orbital plane where the reprocessing column density is highest, while X-ray emission is suppressed in the orbital plane for the same reason. The pericenter is along the $(0^{\circ},90^{\circ})$ ray.
@@
In Figs. 911 we show isotropic equivalent Mollweide skymaps of the Optical + UV (i-UVW2 bands) and the X-ray (Hard+Soft bands) luminosities at t=0.40, 0.79,t=0.40,\,0.79, and $2.08\,t_{\rm fb}$, respectively. Figure 12 also presents ratio maps (Optical+UV/X-ray and Optical+UV/He) at $t=0.79\,t_{\rm fb}$. Near peak luminosity, the Optical+UV emission is strongest along the disrupted star’s orbital plane, consistent with reprocessing of an ionizing continuum in dense debris, whereas the X-ray bands peak at higher inclinations off the plane, where the reprocessing column density is lower (see especially Figs. 1012). At $t=2.08\,t_{\rm fb}$ a localized X-ray bright spot appears (Fig. 11), density and temperature slices indicate it arises from hot gas breaking out of the inner disk (Figs. 1415).
We further quantify the observable features of this simulated TDE by plotting sky-averaged multiband light curves, in Fig. 13. The sky-averaged light curves exhibit a clear X‑ray flash, peaking at $\approx 2.5$ days after disruption (Fig. 13). This result appears consistent with earlier predictions of similar X-ray flashes in supermassive black hole TDEs from Steinberg and Stone (2024), which appear to have been observed in the (supermassive) TDE AT 2022dsb (Malyali et al., 2024). We note the presence of a brief, earlier spike at $t\approx 0.6\ \text{days}$. As stated in the caption of Fig. 13, we regard this earliest spike as likely being a numerical artifact associated with under-resolution of the returning stream tip near pericenter at early times, when its density is low. Similar early flashes of radiation from the stream tip were seen in preliminary, under-resolved runs in Martire et al. (2025). In contrast to Steinberg and Stone (2024), we have made self-consistent predictions for the time-dependent X-ray luminosity in this paper using our novel multigroup FLD transfer module; with grey simulations, spectral predictions can only be made in post-processing, which may not be self-consistent in situations like TDE circularization where (frequency-dependent) radiation pressure dominates over purely hydrodynamical forces.
@@
Figure 13: Light curves for the different energy groups. Most radiation emerges in UV wavelengths; primarily in the unobservable EUV bands (H and He). UV bands (including Swift UVW2) reach peaks or plateaus near $t\approx 1t_{\rm fb}$, while optical wavelengths monotonically rise for the duration of the simulation. The earliest X-ray flash (at $t\approx 0.6~{\rm day}$) is likely an unphysical artifact associated with under-resolving the stream tip due to its low density, as it returns to pericenter at early times (Martire et al., 2025), but the broader X-ray flare from $\approx 1$ to $\approx 2.5$ days is more likely physical, and agrees qualitatively with both the simulation of Steinberg and Stone (2024) and observations of the TDE AT 2022dsb (Malyali et al., 2024). This paper’s multigroup simulations are the first self-consistent prediction of an early X-ray flash in an end-to-end simulation, suggesting that these flares may offer observational insight into both circularization mechanisms and viewing angles.
@@ VI Conclusions
Our implementation of multigroup FLD into rich expands the capabilities of moving-mesh RHD at the cost of greater complexity and computational expense. The multi-step RHD algorithm is described in more detail in §IV, but in brief, the code first updates the hydrodynamic variables, then moves the Voronoi mesh to follow the bulk fluid flow, and then updates the FLD radiation equation [^3] This overview is slightly oversimplified; for example, radiation advection occurs during the hydrodynamics sub-step.. Finally, radiation forces update the gas momentum and energy variables. To mitigate the greater computational cost of performing the RHD in a multigroup way, we have implemented a novel scheme to accelerate convergence in optically thick cells, with minimal loss of accuracy, by limiting the absorption coefficients (as can be seen in Fig. 7).
@@
As a final test for multigroup rich, we have gone beyond test problems with analytic solutions and have run the first end-to-end three dimensional simulation of a TDE in multigroup RHD. Our simulation follows the disruption of a lower main sequence star by a $M=10^{4}M_{\odot}$ IMBH, using astrophysical parameters identical to the grey RHD simulations recently presented in Martire et al. (2025). This simulation produced multi-band light curves across 10 different frequency groups, and runs for a little over 2 fallback times. Although we will defer more detailed explorations of multigroup TDE simulations for future work (Giron et al. in prep), this proof-of-principle simulation is already quite interesting. The bulk hydrodynamics and bolometric light curve appear broadly consistent with the slow circularization seen in Martire et al. (2025), but the emergent radiation field exhibits a complex angular dependence and frequency evolution. Optical and near-UV bands rise monotonically but never make up more than $\sim 1\%$ of the bolometric emission, which is released primarily in the unobservable EUV (although we note that the far-UV Swift UVW2 band does eventually reach $\sim 10\%$ of the bolometric luminosity).